OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (kirbykirby):

The trick is to go from the inside out. If \( f(x)=-4x+7\), then: \[f(5) =-4(5)+7\\ f(t)=-4t+7\\ f(y)=-4y+7\\ f(x+5)=-4(x+1)+7\\ f(g(x))=-4(g(x))+7\] Do you see the logic? The thing inside \(f(.)\) is is just replaced in the expression wherever you see the variable \(x\)

OpenStudy (kirbykirby):

sorry the 4th line should read \(f(x+5)=-4(x+5)+7\)

OpenStudy (anonymous):

OH , okay thank you :)

OpenStudy (kirbykirby):

:)